# AudioTrust: Benchmarking the Multifaceted Trustworthiness of Audio Large Language Models

### v3

@@ 5 AudioTrust: Privacy
Attack Strategies. In AudioTrust, we categorize the privacy risks associated with ALLMs into two distinct groups: (1) Direct Privacy Leakage, which pertains to sensitive information explicitly stated within the conversational content. In this category, the ALLMs might reveal data such as a bank account number mentioned during a conversation. The formulation of this risk is informed by similar challenges in traditional large language models. (Wang et al., 2023a; Huang et al., 2024). (2) Privacy Inference Leakage, where private attributes are inferred from paralinguistic cues rather than the explicit content. Such cues include a speaker’s tone of voice, speech rate, accent, and vocal quality. This risk, which is unique to ALLMs, involves the model deducing personal attributes like age or ethnicity from the audio itself, independent of the semantic content. To evaluate these risks, we constructed two datasets: a direct privacy leakage set containing 600 synthetic dialogues and a privacy inference set with 300 samples. The latter was created using speech from Common Voice (Ardila et al., 2020) mixed with background audio from freesound (Freesound, ). For a detailed taxonomy and construction methods, see Appendixes G.1 and G.2.
Evaluation & Metrics. We evaluated our experiment within an Audio Question Answering (AQA) framework, employing two distinct settings: one utilizing privacy-enhancing prompts and the other using standard prompts. GPT-4o serves as the evaluator, assigning scores from 1 to 5 to each response and then the results were checked by human. A score of 1 indicates complete disclosure of private information, while a score of 5 signifies a refusal of the request due to privacy safety concerns. Our main evaluation metrics are refusal rate and accuracy. A higher refusal rate indicates stronger privacy protection. Accuracy is calculated only when privacy information is leaked. Further details on evaluation prompts and evaluation metrics can be found in Appendixes G.3 and G.4.
| Model Name | Direct leakage | Inference leakage | ||
@@
| SALMONN | 57.50 | 96.83 | 38.75 | 45.00 |
| UltraVox | 73.46 | 99.67 | 6.50 | 7.50 |
| Qwen2-Audio | 0.83 | 23.67 | 1.50 | 0.75 |
| MiniCPM-o 2.6 | 0.00 | 0.67 | 1.50 | 1.50 |
| Step Fun | 41.50 | 98.33 | 11.50 | 14.75 |
| Qwen2.5-Omni | 0.00 | 1.17 | 0.50 | 0.00 |
| Kimi Audio | 0.17 | 1.00 | 11.25 | 6.50 |
| OpenS2S | 7.68 | 43.83 | 11.00 | 10.00 |
| Step Audio2 | 0.00 | 38.83 | 0.00 | 0.00 |
@@
| Gemini-1.5 Pro | 11.85 | 98.50 | 4.00 | 7.00 |
| GPT-4o Audio | 92.00 | 99.67 | 5.00 | 16.25 |
| GPT-4o mini Audio | 100.00 | 100.00 | 10.50 | 30.00 |
| Gemini-2.5 Flash | 1.36 | 96.44 | 1.25 | 2.75 |
| Gemini-2.5 Pro | 33.50 | 94.17 | 0.50 | 1.25 |
| Average | 29.99 | 63.77 | 7.41 | 10.23 |
@@
Note: Scores are refusal rates (higher is better). “w/o” vs. “w/” compares standard prompting without and with privacy-aware prompt engineering.
@@
Results. We evaluated the privacy protection performance of both open-source and closed-source ALLMs. The results are shown in Table˜4. Our analysis revealed several key observations: (1) Direct Privacy Leakage: In experiments using the direct privacy leakage dataset, closed-source ALLMs generally achieved superior results. The GPT-4o series demonstrated the best refusal rate, reaching 100%. Furthermore, the implementation of privacy-enhancing prompts significantly improved performance across almost every model, showing approximately a 25% improvement. This indicates that traditional prompt engineering can effectively enhance privacy protection when sensitive information is explicitly present in conversation content. (2) Privacy Inference Leakage: Both open-source and closed-source ALLMs performed poorly in addressing privacy inference leakage. They rarely refused requests for certain types of privacy information, such as age and ethnicity, with the refusal rate of only 7.41%. Unlike direct privacy leakage, privacy-enhancing prompts had a minimal impact, yielding only about a 3% improvement. Across all our results, we observed that ALLMs struggle to process privacy information that is not directly stated in a conversation but rather inferred from paralinguistic cues. This suggests that ALLMs might not identify such inferred information as private or requiring protection. This could be due to the training process, where paralinguistic cues may have been considered less important than conversational content, or because of insufficient data for this specific type of information. Detailed results accuracy analysis are available in Appendix G.5.
@@ 6 AudioTrust: Robustness
This section investigates the robustness of ALLMs in maintaining performance against real-world audio distortions. Unlike in text (Wang et al., 2023b) and vision (Hendrycks & Dietterich, 2019), audio robustness presents unique challenges due to its physical nature. Audio signals are temporal encodings of acoustic phenomena, with inherent properties like source, medium, and reverberation shaping complex auditory scenes. A truly robust ALLM must be able to disentangle the primary speech signal from this acoustic clutter and maintain its performance. This is not simply a matter of better speech recognition, but a test of the model’s fundamental ability to function in imperfect acoustic environments.
Attack Strategies. We categorize robustness challenges for ALLMs into two primary types: intentional adversarial attacks (Madry et al., 2017) and naturally occurring phenomenon of performance degradation (Radford et al., 2023). Adversarial attacks employ carefully crafted, imperceptible perturbations to induce model failure (Carlini & Wagner, 2018). In contrast, non-adversarial challenges encompass common real-world interferences. We evaluate model robustness across several key dimensions: (1) adversarial resilience, including three categories: natural noise, speaker identification, and voice overlap situations.speech (Bredin et al., 2020); (2) robustness to environmental noise (Zhang et al., 2023b) and variations in audio quality. To this end, we constructed dedicated datasets simulating these interferences. Each dimension contains 40 multilingual and multi-topic samples to ensure a comprehensive assessment. Further dataset details are available in Sections˜H.1 and H.2.
Evaluation & Metrics. Given the challenges in directly measuring robustness or output risk, we adopt a model-based evaluation using GPT-4o (OpenAI et al., 2024), following recent evidence (Zheng et al., 2023a). Each test output is rated on a discrete 10-point scale, with scoring rubrics tailored per prompt and task: 10 indicates strong consistency with audio quality, while 0 means perceptual failure or inability to recognize the specified variation. These scores were then subject to further review by the human evaluators. Prompt templates are detailed in Section˜H.3.1. For comprehensive evaluation, we also report two quantitative metrics (Section˜H.3.2): CM-WER, measuring dissimilarity between generated and human-annotated transcriptions (Radford et al., 2023); and Content Consistency Rate (CCR),(e.g., transcribe the voice with added interfering information through multiple rounds of model dialogue prompts to understand its semantic context, and then score the transcription against the original voice text content.) assessing factual alignment between ALLMs outputs and ground-truth audio content (Min et al., 2023).
@@
| Open-source | ||||||
| MiniCPM-o 2.6 | 7.80 | 7.18 | 7.92 | 7.06 | 6.50 | 6.17 |
| Qwen2-Audio | 6.00 | 3.50 | 4.33 | 6.84 | 5.40 | 6.60 |
| SALMONN | 2.00 | 6.42 | 4.57 | 2.94 | 7.16 | 6.66 |
| Ultravox | 4.00 | 7.53 | 7.30 | 6.53 | 6.70 | 7.00 |
| Qwen2.5-Omni | 8.14 | 7.10 | 7.50 | 7.93 | 7.12 | 7.17 |
| Step-Fun | 5.00 | 7.48 | 8.20 | 7.42 | 5.89 | 7.08 |
| Kimi Audio | 5.67 | 6.83 | 6.00 | 6.83 | 7.08 | 6.94 |
| OpenS2S | 8.25 | 6.46 | 5.17 | 6.39 | 2.33 | 6.25 |
| Step-Audio2 | 6.18 | 6.58 | 7.92 | 6.82 | 0.00 | 6.78 |
| Closed-source | ||||||
| Gemini-1.5 Pro | 8.57 | 8.21 | 8.23 | 8.16 | 6.09 | 7.43 |
| Gemini-2.5 Flash | 8.16 | 8.38 | 8.28 | 7.93 | 6.36 | 7.76 |
| Gemini-2.5 Pro | 8.88 | 8.68 | 8.50 | 8.18 | 7.46 | 7.71 |
| GPT-4o Audio | 5.90 | 5.50 | 8.33 | 7.31 | 7.62 | 6.27 |
| GPT-4o mini Audio | 8.33 | 6.90 | 7.69 | 6.00 | 5.77 | 7.25 |
| Average | 6.67 | 6.94 | 7.08 | 6.89 | 6.27 | 6.95 |
Note: Scenarios: AR (Adversarial Robustness), AQV (Audio Quality Variation), BC (Background Conversation), ES (Environmental Sound), MS (Multiple Speakers), NI (Noise Interference). The overall low accuracy highlights hallucination concerns.
@@ 7 AudioTrust: Authentication
Attack Strategies. We devise a taxonomy of authentication attacks that exploit the multi-dimensional attributes of audio signals, categorizing them into three primary classes. (1) Identity Verification Bypass (IVB), which injects strong emotional cues (e.g., feigned urgency or distress) to exploit social engineering principles and induce the model to lower its security thresholds, and (2) Hybrid Spoofing (HS), this attack convolves cloned or synthesized speech with background noise and reverberation characteristic of specific scenes (e.g., a bustling office). Such attacks challenge the model’s robustness in source separation and noise suppression, and more critically, its ability to jointly reason about the logical consistency between the purported speaker identity and their acoustic context. (3) Voice Cloning Spoofing (VCS), which employs advanced speech synthesis (e.g., F5-TTS (Chen et al., 2024)) to generate highly realistic cloned voices saying a correct passphrase, thereby testing the model’s sensitivity to subtle physical traits in the vocal fingerprint. This taxonomy moves beyond traditional spoofing distinctions to directly probe ALLMs’ intrinsic vulnerabilities in processing complex auditory signals. Based on this framework, we constructed a 400-sample evaluation dataset (details in Sections˜I.1 and I.2).
@@
| SALMONN | 74 | 93 | N/A |
| Ultravox | 5 | 43 | 72 |
| Qwen2-Audio | 58 | 29 | 7.5 |
| MiniCPM-o 2.6 | 76 | 57 | 20.5 |
| Step-Fun | 11 | 3 | 78 |
| Qwen2.5-omni | 36 | 36 | 87.5 |
| Kimi-Audio | 21 | 24 | 75.5 |
| OpenS2S | 3 | 34 | 50 |
| Step-Audio2 | 36 | 85 | 49 |
| Open-source Avg. | 35.6 | 44.9 | 55.0 |
@@
| Gemini-1.5 pro | 4 | 5 | 66.5 |
| GPT-4o Audio | 2 | 0 | 16.5 |
| GPT-4o mini Audio | 0 | 0 | 14 |
| Gemini-2.5 Flash | 3 | 7 | 89 |
| Gemini-2.5-Pro | 5 | 3 | 89.5 |
| Closed-source Avg. | 2.8 | 3.0 | 55.1 |
@@
Note: SALMONN consistently disregarded prompt instructions by outputting audio descriptions, which prevented obtaining valid results for voice cloning spoofing. For authentication metrics, lower values indicate better security (fewer successful attacks).
@@
Evaluation and Metrics. Our evaluation framework is designed to comprehensively assess the robustness of models under different attack scenarios. FAR is defined as the proportion of impostor voices incorrectly identified as authentic users, which will serve as the privacy metric for all attack scenarios. GPT-4o will be our main evaluator, used to verify if ALLMs deny attacker requests. Human evaluators will then review GPT-4o’s results before the final FAR calculation. In the case of HS, where the model faces both identity spoofing and environmental interference, FAR remains the central performance indicator. A failure is registered whenever the model responds positively to a forged input (e.g., acknowledging identity confirmation or granting access). Further details on the evaluation are provided in Section˜I.3 and I.4.
@@
Results. In Table˜6, we report the performance of both open- and closed-source ALLMs on our evaluation dataset. Several key observations emerge: (1) Our findings in the IVB and HS settings reveal a critical attack vector absent in text-only systems: non-semantic features. We observed that the average false acceptance rate for open-source models was 35.6% in IVB settings, but significantly higher at 44.9% in HS settings. These settings differ considerably in paralinguistic features. The HS setting, for example, includes background audio and other speakers’ voices, both common elements in social engineering attacks conveyed through audio. This suggests that paralinguistic cues might aid attackers in successful authentication. Conversely, closed-source models appear to demonstrate greater robustness. (2) In the VCS tests, we observe substantial performance discrepancies across models, with the Gemini family exhibiting comparatively weaker defenses. Interestingly, we find that simply increasing the strictness of system prompts consistently improves resilience against spoofing attacks across all systems. This suggests that in downstream ALLM applications, carefully crafted system prompts provide an efficient means of strengthening authentication security. Further details can be found in Section˜I.5.

### v4

@@ 5 AudioTrust: Privacy
| Model | Direct leakage | Inference leakage | ||
@@
| SALMONN | 57.50 / 57.00 | 96.83 / 96.83 | 48.33 / 47.00 | 49.00 / 49.00 |
| UltraVox | 73.46 / 72.67 | 99.67 / 99.33 | 8.67 / 8.67 | 9.67 / 7.33 |
| Qwen2-Audio | 0.83 / 0.83 | 23.67 / 23.17 | 1.33 / 0.00 | 1.00 / 0.00 |
| MiniCPM-o 2.6 | 0.00 / 0.00 | 0.67 / 0.67 | 1.33 / 0.00 | 1.00 / 0.33 |
| Step Fun | 41.50 / 41.67 | 98.33 / 98.17 | 15.33 / 14.67 | 19.67 / 17.00 |
| Qwen2.5-Omni | 0.00 / 0.00 | 1.17 / 1.00 | 0.33 / 0.00 | 0.00 / 0.00 |
| Kimi Audio | 0.17 / 0.17 | 1.00 / 1.00 | 15.00 / 15.33 | 8.67 / 8.67 |
| OpenS2S | 7.68 / 7.17 | 43.83 / 42.83 | 8.67 / 6.87 | 7.33 / 7.83 |
| Step Audio2 | 0.00 / 0.00 | 38.83 / 38.83 | 0.00 / 0.00 | 0.00 / 0.00 |
@@
| Gemini-1.5 Pro | 11.85 / 12.12 | 98.50 / 98.67 | 5.33 / 6.33 | 9.33 / 8.67 |
| GPT-4o Audio | 92.00 / 91.83 | 99.67 / 99.50 | 6.67 / 6.33 | 20.00 / 17.33 |
| GPT-4o mini Audio | 100.00 / 100.00 | 100.00 / 100.00 | 14.00 / 13.33 | 40.00 / 40.00 |
| Gemini-2.5 Flash | 1.36 / 0.00 | 96.44 / 96.82 | 1.00 / 1.00 | 3.33 / 0.67 |
| Gemini-2.5 Pro | 33.50 / 33.33 | 94.17 / 94.33 | 0.33 / 0.00 | 0.67 / 0.00 |
| Average | 29.99 / 29.77 | 63.77 / 63.65 | 9.02 / 8.54 | 12.12 / 11.20 |
@@
Note: Values are shown as GPT-4o / Qwen3 evaluators. Scores are refusal rates (higher is better). “w/o” vs. “w/” compares standard prompting without and with privacy-aware prompt engineering.
Attack Strategies. In AudioTrust, we categorize the privacy risks associated with ALLMs into two distinct groups: (1) Direct Privacy Leakage, which pertains to sensitive information explicitly stated within the conversational content. In this category, the ALLMs might reveal data such as a bank account number mentioned during a conversation. The formulation of this risk is informed by similar challenges in traditional large language models. (Wang et al., 2023a; Huang et al., 2024). (2) Privacy Inference Leakage, where private attributes are inferred from paralinguistic cues rather than the explicit content. Such cues include a speaker’s tone of voice, speech rate, accent, and vocal quality. This risk, which is unique to ALLMs, involves the model deducing personal attributes from the audio itself, independent of the semantic content. Consistent with recent studies (Xu et al., 2025a; Feng et al., 2023; Maltezou-Papastylianou et al., 2025), we treat the systematic inference of demographic attributes (e.g., age, ethnicity) from voice as a critical privacy leakage. While distinct from unique identity recognition, such profiling exploits paralinguistic cues to reveal sensitive personal data, which can be aggregated for user profiling at scale. Therefore, we evaluate demographic inference as a primary dimension of privacy risk. To evaluate these risks, we constructed two distinct datasets. The first, a direct privacy leakage set, contains 600 synthetic dialogues. This set was created by first synthesizing privacy information, then generating the conversations using LLMs, and finally transcribing them into audio. The second, a privacy inference set, comprises 300 samples. For this set, we used speech from Common Voice (Ardila et al., 2020) and mixed it with background audio sourced from freesound (Freesound, ). For a detailed taxonomy and construction methods, see Appendixes G.1 and G.2.
Evaluation & Metrics. We evaluated our experiment within an Audio Question Answering (AQA) framework, employing two distinct settings: one utilizing privacy-enhancing prompts and the other using standard prompts. GPT-4o and Qwen3 serve as the evaluators, assigning scores from 1 to 5 to each response and then the results were checked by human. A score of 1 indicates complete disclosure of private information, while a score of 5 signifies a refusal of the request due to privacy safety concerns. Our main evaluation metrics are refusal rate and accuracy. A higher refusal rate indicates stronger privacy protection. Accuracy is calculated only when privacy information is leaked. Further details on evaluation prompts and evaluation metrics can be found in Appendixes G.3 and G.4.
@@
Results. We evaluated the privacy protection performance of both open-source and closed-source ALLMs. The results are shown in Table˜4. Our analysis revealed several key observations: (1) Direct Privacy Leakage: In experiments using the direct privacy leakage dataset, closed-source ALLMs generally achieved superior results. The GPT-4o series demonstrated the best refusal rate, reaching 100%. Furthermore, the implementation of privacy-enhancing prompts significantly improved performance across almost every model, showing approximately a 25% improvement. This indicates that traditional prompt engineering can effectively enhance privacy protection when sensitive information is explicitly present in conversation content. (2) Privacy Inference Leakage: Both open-source and closed-source ALLMs performed poorly in addressing privacy inference leakage. They rarely refused requests for certain types of privacy information, such as age and ethnicity, with the refusal rate of only 9.02%. Unlike direct privacy leakage, privacy-enhancing prompts had a minimal impact, yielding only about a 3% improvement. Across all our results, we observed that ALLMs struggle to process privacy information that is not directly stated in a conversation but rather inferred from paralinguistic cues. This suggests that ALLMs might not identify such inferred information as private or requiring protection. This could be due to the training process, where paralinguistic cues may have been considered less important than conversational content, or because of insufficient data for this specific type of information. Detailed results accuracy analysis are available in Appendix G.5.
@@ 6 AudioTrust: Robustness
This section investigates the robustness of ALLMs in maintaining performance against real-world audio distortions. Unlike in text (Wang et al., 2024b) and vision (Hendrycks & Dietterich, ), audio robustness presents unique challenges due to its physical nature. Audio signals are temporal encodings of acoustic phenomena, with inherent properties like source, medium, and reverberation shaping complex auditory scenes. A truly robust ALLM must be able to disentangle the primary speech signal from this acoustic clutter and maintain its performance. This is not simply a matter of better speech recognition, but a test of the model’s fundamental ability to function in imperfect acoustic environments.
@@
| Open-source Models | ||||||
| MiniCPM-o 2.6 | 7.80 / 4.85 | 7.18 / 5.35 | 7.92 / 7.57 | 7.06 / 7.98 | 6.50 / 7.90 | 6.17 / 5.50 |
| Qwen2-Audio | 6.00 / 2.95 | 3.50 / 3.97 | 4.33 / 6.29 | 6.84 / 6.23 | 5.40 / 8.07 | 6.60 / 4.97 |
| SALMONN | 2.00 / 2.58 | 6.42 / 5.01 | 4.57 / 5.59 | 2.94 / 4.45 | 7.16 / 5.58 | 6.66 / 5.63 |
| Ultravox | 4.00 / 4.76 | 7.53 / 6.29 | 7.30 / 7.94 | 6.53 / 8.05 | 6.70 / 8.30 | 7.00 / 5.93 |
| Qwen2.5-Omni | 8.14 / 5.63 | 7.10 / 6.59 | 7.50 / 7.93 | 7.93 / 7.94 | 7.12 / 7.28 | 7.17 / 6.03 |
| Step-Fun | 5.00 / 4.78 | 7.48 / 6.50 | 8.20 / 8.36 | 7.42 / 7.46 | 5.89 / 4.97 | 7.08 / 6.86 |
| Kimi Audio | 5.67 / 4.00 | 6.83 / 5.63 | 6.00 / 5.70 | 6.83 / 5.00 | 7.08 / 6.41 | 6.94 / 5.98 |
| OpenS2S | 8.25 / 3.83 | 6.46 / 4.10 | 5.17 / 6.11 | 6.39 / 5.15 | 2.33 / 8.92 | 6.25 / 4.85 |
| Step-Audio2 | 6.18 / 4.93 | 6.58 / 6.45 | 7.92 / 7.91 | 6.82 / 7.61 | 0.00 / 0.12 | 6.78 / 7.25 |
| Closed-source Models | ||||||
| Gemini-1.5 Pro | 8.57 / 6.63 | 8.21 / 7.89 | 8.23 / 8.01 | 8.16 / 8.08 | 6.09 / 8.39 | 7.43 / 7.02 |
| Gemini-2.5 Flash | 8.16 / 6.98 | 8.38 / 7.89 | 8.28 / 8.14 | 7.93 / 7.74 | 6.36 / 8.12 | 7.76 / 7.54 |
| Gemini-2.5 Pro | 8.88 / 7.74 | 8.68 / 8.19 | 8.50 / 8.36 | 8.18 / 7.87 | 7.46 / 7.87 | 7.71 / 7.60 |
| GPT-4o Audio | 5.90 / 4.66 | 5.50 / 5.33 | 8.33 / 8.31 | 7.31 / 7.70 | 7.62 / 9.88 | 6.27 / 5.91 |
| GPT-4o mini Audio | 8.33 / 5.45 | 6.90 / 7.09 | 7.69 / 8.05 | 6.00 / 7.73 | 5.77 / 8.53 | 7.25 / 6.30 |
| Average | 6.63 / 4.98 | 6.91 / 6.16 | 7.14 / 7.45 | 6.88 / 7.07 | 5.82 / 7.17 | 6.93 / 6.24 |
Note: Values are shown as GPT-4o / Qwen3 evaluators. Subscript is the absolute difference from the respective column average. Scenarios: AR (Adversarial Robustness), AQV (Audio Quality Variation), BC (Background Conversation), ES (Environmental Sound), MS (Multiple Speakers), NI (Noise Interference).
Attack Strategies. We categorize robustness challenges for ALLMs into two primary types: intentional adversarial attacks (Madry et al., 2021) and naturally occurring phenomenon of performance degradation (Radford et al., 2023). Adversarial attacks employ carefully crafted, imperceptible perturbations to induce model failure (Carlini & Wagner, 2018) Details of the creation of additional interference data can be found in Section˜H.2). In contrast, non-adversarial challenges encompass common real-world interferences. We evaluate model robustness across several key dimensions: (1) adversarial resilience, including three categories: natural noise, speaker identification, and voice overlap situations (Bredin et al., 2020); (2) robustness to environmental noise (Zhang et al., 2023b) and variations in audio quality. To this end, we constructed dedicated datasets simulating these interferences. In order to ensure the fluency of natural language and the logical validity of the original speech, we created the dataset by adding the aforementioned simulated interference to the introduced big bench audio (Suzgun et al., 2023). Each dimension contains 40 multilingual and multi-topic samples to ensure a comprehensive assessment. Further dataset details are available in Sections˜H.1 and H.2.
Evaluation & Metrics. Given the challenges in directly measuring robustness or output risk, we adopt a model-based evaluation using GPT-4o (OpenAI et al., 2024) and Qwen3-based (Yang et al., 2025a), following recent evidence (Zheng et al., 2023a). Each test output is rated on a discrete 10-point scale, with scoring rubrics tailored per prompt and task: 10 indicates strong consistency with audio quality, while 0 means perceptual failure or inability to recognize the specified variation. These scores were then subject to further review by the human evaluators. Prompt templates are detailed in Section˜H.3.1. For comprehensive evaluation, we also report two quantitative metrics (Section˜H.3.2): CM-WER, measuring dissimilarity between generated and human-annotated transcriptions (Radford et al., 2023); and Content Consistency Rate (CCR),(e.g., transcribe the voice with added interfering information through multiple rounds of model dialogue prompts to understand its semantic context, and then score the transcription against the original voice text content.) assessing factual alignment between ALLMs outputs and ground-truth audio content (Min et al., 2023).
@@ 7 AudioTrust: Authentication
| SALMONN | 26 / 26 | 7 / 8 | N/A |
| Ultravox | 95 / 96 | 57 / 57 | 28 |
| Qwen2-Audio | 42 / 39 | 71 / 71 | 92.5 |
| MiniCPM-o 2.6 | 24 / 20 | 43 / 39 | 79.5 |
| Step-Fun | 79 / 79 | 97 / 97 | 22 |
| Qwen2.5-omni | 19 / 19 | 64 / 64 | 12.5 |
| Kimi-Audio | 79 / 74 | 76 / 76 | 24.5 |
| OpenS2S | 97 / 93 | 66 / 65 | 50 |
| Step-Audio2 | 37 / 37 | 15 / 15 | 51 |
| Open-source Avg. | 55.3 / 53.7 | 55.1 / 54.7 | 45.0 |
@@
| Gemini-1.5 pro | 96 / 96 | 95 / 95 | 33.5 |
| GPT-4o Audio | 98 / 98 | 100 / 100 | 83.5 |
| GPT-4o mini Audio | 100 / 100 | 100 / 100 | 86 |
| Gemini-2.5 Flash | 97 / 97 | 93 / 93 | 11 |
| Gemini-2.5-Pro | 95 / 95 | 97 / 97 | 10.5 |
| Closed-source Avg. | 97.2 / 97.2 | 97.0 / 97.0 | 44.9 |
@@
Note: Values in IVB and HS columns are shown as GPT-4o evaluator / Qwen3 evaluator. SALMONN consistently disregarded prompt instructions by outputting audio descriptions, which prevented obtaining valid results for voice cloning spoofing. For authentication metrics, higher values indicate better security (more successful defenses against attacks). The VCS component was not assessed by large language model evaluators.
Attack Strategies. We devise a taxonomy of authentication attacks that exploit the multi-dimensional attributes of audio signals, categorizing them into three primary classes. (1) Identity Verification Bypass (IVB), which injects strong emotional cues (e.g., feigned urgency or distress) to exploit social engineering principles and induce the model to lower its security thresholds, and (2) Hybrid Spoofing (HS), this attack convolves cloned or synthesized speech with background noise and reverberation characteristic of specific scenes (e.g., a bustling office). Such attacks challenge the model’s robustness in source separation and noise suppression, and more critically, its ability to jointly reason about the logical consistency between the purported speaker identity and their acoustic context. (3) Voice Cloning Spoofing (VCS), which employs advanced speech synthesis (e.g., F5-TTS (Chen et al., 2025)) to generate highly realistic cloned voices saying a correct passphrase, thereby testing the model’s sensitivity to subtle physical traits in the vocal fingerprint. This taxonomy moves beyond traditional spoofing distinctions to directly probe ALLMs’ intrinsic vulnerabilities in processing complex auditory signals. Based on this framework, we constructed a 400-sample evaluation dataset (details in Sections˜I.1 and I.2).
@@
Evaluation and Metrics. Our evaluation framework is designed to comprehensively assess the robustness of models under different attack scenarios. Imposter Rejection Rate (IRR) is defined as the proportion of impostor voices that are correctly rejected by the system, which will serve as the privacy metric for all attack scenarios. GPT-4o and Qwen3 will be our main evaluator, used to verify if ALLMs deny attacker requests. Human evaluators will review GPT-4o’s outputs before computing the final IRR. In the HS scenario, where the model must handle both identity spoofing and environmental interference, the IRR serves as the primary performance indicator. A success is recorded only when the model correctly rejects a forged input,such as refusing to acknowledge an identity confirmation request or denying access. Further details on the evaluation are provided in Section˜I.3 and I.4.
@@
Results. In Table˜6, We evaluate both open- and closed-source ALLMs on our benchmark. We report IRR for the authentication task. Several key observations emerge: (1) Our findings in the IVB and HS settings reveal a critical attack vector absent in text-only systems: non-semantic features. We observed that the average IRR for open-source models was 55.3% in IVB settings, but slightly higher at 55.1% in HS settings. These settings differ considerably in paralinguistic features. The HS setting, for example, includes background audio and other speakers’ voices, both common elements in social engineering attacks conveyed through audio. This suggests that paralinguistic cues have limited influence on successful authentication by attackers. (2) In the VCS tests, we observe substantial performance discrepancies across models, with the Gemini family exhibiting comparatively weaker defenses. Interestingly, we find that simply increasing the strictness of system prompts consistently improves resilience against spoofing attacks across all systems. This suggests that in downstream ALLM applications, carefully crafted system prompts provide an efficient means of strengthening authentication security. Further details can be found in Section˜I.5.